OpenStudy (anonymous):

Identify the slope and y-intercept of the inequality 12x + 3y > 15. slope = 4, y-intercept = 5 slope = −4, y-intercept = 15 slope = negative 1 over 4, y-intercept = 5 slope = −4, y-intercept = 5

OpenStudy (raden):

if given a straight line ax+by=c, the slope or gradien (m) = -a/b to find y-intercept, take x=0 and next solve for y

OpenStudy (anonymous):

Im just trying to get finished with this exam, I already passed this course, my mom made me re-take it on florida virtual school

OpenStudy (raden):

from ur equation, given 12x + 3y > 15 it mean a=12, b=3 so, just plug the value of a and b to the formula like i said before show what u get